# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ Abstract
The stability of ecological systems is a fundamental concept in ecology, which offers profound insights into species coexistence, biodiversity, and community persistence. In this article, we provide a systematic and comprehensive review on the theoretical frameworks for analyzing the stability of ecological systems. Notably, we survey various stability notions, including linear stability, sign stability, diagonal stability, D-stability, total stability, structural stability, and higher-order stability. For each of these stability notions, we examine necessary or sufficient conditions for achieving such stability and demonstrate the intricate interplay of these conditions on the network structures of ecological systems. Finally, we explore the future prospects of these stability notions.
@@
{AMS}
92D25, 92D40, 93C15, 93D05, 93D20
Contents
1  Introduction 1
2  Preliminaries 2
2.1  Definitions of Lyapunov stability 2.1
2.2  Lyapunov’s indirect method 2.2
2.3  Lyapunov’s direct method 2.3
3  Linear stability 3
3.1  Model-implicit approaches 3.1
3.1.1  May’s classical result 3.1.1
3.1.2  Impacts of interaction types and correlation 3.1.2
3.1.3  Impacts of degree heterogeneity 3.1.3
3.1.4  Impacts of self-regulation 3.1.4
3.1.5  Impacts of modularity 3.1.5
3.1.6  Impacts of dispersal 3.1.6
3.1.7  Impacts of time delay 3.1.7
3.1.8  Limitations of model-implicit approaches 3.1.8
3.2  Model-explicit approaches 3.2
3.2.1  Impacts of equilibria 3.2.1
3.2.2  Impacts of extinction 3.2.2
3.2.3  Impacts of time delay 3.2.3
3.2.4  Impacts of stochastic noises 3.2.4
3.2.5  Impacts of evolved system size 3.2.5
3.2.6  Applications to real ecological systems 3.2.6
4  Sign stability 4
4.1  Characterizations of sign stable matrices 4.1
4.2  Applications to ecological systems 4.2
5  Diagonal stability 5
5.1  Lyapunov stability of the GLV model 5.1
5.2  Characterizations of diagonally stable matrices 5.2
5.2.1  Negative feedback cyclic structure 5.2.1
5.2.2  Cactus structure 5.2.2
5.2.3  Rank-one structure 5.2.3
6  D-stability 6
6.1  Characterizations of D-stable matrices 6.1
6.2  Results with the GLV model 6.2
7  Total stability 7
8  Structural stability 8
8.1  Mathematical definition 8.1
8.2  Structural stability metrics 8.2
9  Higher-order stability 9
9.1  GLV model with higher-order interactions 9.1
9.2  Two potential approaches 9.2
9.2.1  Lyapunov approach 9.2.1
9.2.2  Tensor decomposition approach 9.2.2
9.3  Implicit higher-order interactions 9.3
9.3.1  Impacts of cascade structure 9.3.1
9.3.2  Impacts of mutualistic interactions 9.3.2
9.3.3  Impacts of warming temperature 9.3.3
9.3.4  Impacts of dynamic switching 9.3.4
9.3.5  Impacts of periodic environments 9.3.5
9.3.6  Impacts of consumption threshold 9.3.6
9.3.7  Impacts of delayed age structure 9.3.7
9.3.8  A case study of larch moth interactions 9.3.8
10  Discussion 10
11  Conclusion 11
Reference 46
@@ 1 Introduction
Ecological systems are usually subject to continual disturbances or perturbations, and their responses are often characterized quantitatively as stability, a fundamental concept rooted in dynamical systems and control theory [56, 136, 154, 175, 185, 225, 226, 231, 254]. Loosely speaking, an ecological system is stable if its state represented by a vector of its species abundances, do not change too much under small perturbations. Therefore, understanding the stability of ecological systems is particularly important due to its implications for species coexistence, biodiversity, and community persistence [74, 105, 113, 119, 123, 152, 184, 253]. More importantly, these implications can contribute to the development of effective interventions and policies that maintain the overall health of ecological systems [81, 95, 129].
@@
Numerous mathematicians, physicists, and ecologists have explored the stability of ecological systems. In his pioneer work, May linearized an unspecified nonlinear dynamical system around an equilibrium to perform the local linear stability analysis of ecological systems [163, 165, 166]. The local linear stability is completely characterized by the eigenvalues of the Jacobian matrix (of the unspecified dynamical model), which is often called the community matrix in the ecological contexts. If all the eigenvalues of the community matrix have negative real parts, the ecological system is locally stable around the equilibrium. By incorporating random matrix theory, May successfully derived a simple condition to characterize the stability of ecological systems. While there have been intense debates about the relationship between stability and complexity over the past five decades [168, 181, 199, 211, 230, 256, 260, 275], May’s result offers a powerful foundation for analyzing the stability of ecological systems with diverse structural characteristics, such as interaction types [11], degree heterogeneity [270], self-regulation [30], modularity [112], and more.
@@
May’s framework does not need to know the dynamical model of the ecological system. Instead, it focuses on the Jacobian matrix of the unspecified dynamical model. Hence, the framework is model-implicit. In the past, actually many ecological models, such as the generalized Lotka-Volterra (GLV) model (1910) [49, 160], Holling type II model (1959) [2, 235], and MacArthur’s consumer-resource model (1970) [1, 70], have been proposed to study the dynamics of ecological systems. Remarkably, the GLV model is the most commonly used model in the stability analysis of ecological systems due to its simplicity [5, 13, 42, 43, 57, 107, 142, 277]. Most of these findings rely on Lyapunov theory, which allows us to determine the stability of a system without explicitly integrating the differential equation. Lyapunov theory encompasses two fundamental methods. One is based on linearization (e.g., May’s approach), which is often referred to as the Lyapunov indirect method, while the other is called the Lyapunov direct method. The essence of Lyapunov direct method is to construct the so-called Lyapunov function, an “energy-like” scalar function whose time variation can be viewed as the “energy dissipation.” It is not restricted to small perturbations, and in principle can be applied to all dynamical systems. However, we lack a general theory to find a suitable Lyapunov function for an arbitrary dynamical system. Instead, we have to rely on our experience and intuition. Notably, certain diagonal-type Lyapunov functions are useful for the stability analysis of ecological systems described by the GLV model [135].
@@
The feasibility of an ecological system, which concerns whether a specific set of conditions or parameters can support the system over time, can also impact its stability [7, 18, 106, 110, 187, 244]. Intriguingly, the relationship between feasibility and stability is noteworthy when dealing with ecological systems described by the GLV model or MacArthur’s consumer-resource model. It has been proved that under certain conditions, the feasibility of such systems can imply their stability [59, 244]. Moreover, feasibility is closely related to a concept known as structural stability, which reflects the ability of an ecological system to qualitatively maintain its dynamics and overall behavior under small disturbances or perturbations of the system model itself [213]. A great amount of metrics have been proposed to quantify the structural stability of ecological systems through feasibility analysis [110, 213, 220].
Recently, there has been a growing interest in higher-order interactions within ecological systems, which refers to the intricate relationships that extend beyond pairwise interspecies interactions [27, 32, 101, 111, 153], see Fig. 1a, b. Mathematically, these higher-order interactions emerged in an ecological system can be naturally represented as a hypergraph, where its hyperedges can connect an arbitrary number of nodes [44]. The resulting dynamics can be expressed in the form of polynomials [67]. The analysis of hypergraph dynamics often involves the application of tensor theory [65, 67, 197, 247], which deals with multidimensional arrays generalized from vectors and matrices [62, 66, 68]. Lyapunov theory can also be leveraged to analyze the stability of polynomial dynamical systems [3]. Furthermore, the consumer-resource model can be implicitly considered as an instance of higher-order interactions, with consumer species as nodes and resources as hyperedges, see Fig. 1c. Increasing evidence has revealed that higher-order interactions play a significant role in the stability of ecological systems [27, 101, 111].
Table 1: Summary of various stability notions that have been used to study ecological systems.
@@
[9, 11, 30, 36, 38, 96, 100, 109, 112, 126, 146, 163, 165, 198, 203, 229, 244, 249, 269, 270, 271]
@@
[10, 77, 118, 130, 158, 167, 206, 268]
@@
[19, 20, 21, 22, 42, 43, 100, 107, 110, 120, 124, 135, 142, 171, 210, 233, 262, 266]
@@
[24, 72, 90, 100, 132, 213]
@@
[156, 157, 159, 206]
@@
[17, 110, 190, 196, 209, 213, 218, 220, 237, 238, 242]
@@
[3, 6, 4, 18, 27, 51, 58, 59, 62, 78, 80, 84, 87, 98, 101, 111, 137, 227, 234, 248, 276]
Most significantly, the stability of an ecological system is profoundly determined and characterized by its underlying network structure, which captures how species influence one another and how disturbances or perturbations propagate among species. In this article, we provide a systematic and inclusive literature review on the theoretical frameworks for analyzing the stability of ecological systems. In Section 2, we briefly review the fundamentals of Lyapunov theory, which acts as a foundation for the stability analysis of ecological systems. From Section 3 to Section 9, we comprehensively survey various stability notions, including linear stability, sign stability, diagonal stability, D-stability, total stability, structural stability, and higher-order stability, see Table 1. For each of these stability notions, we examine necessary or sufficient conditions that are required to establish such stability and illustrate the complicated relationships between these conditions and the network structures of ecological systems. Finally, we discuss the future prospects of these stability notions in Section 10 and conclude in Section 11.
@@ 2 Preliminaries
In this section, we review the fundamentals of Lyapunov theory, which plays a significant role in the various stability analyses of ecological systems. We adapt most of the notations and results from the comprehensive work of [138, 139, 172, 224, 236].
@@ 2.1 Definitions of Lyapunov stability
Otherwise, ${\bf x}^{*}$ is unstable, see Fig. 2. An equilibrium point ${\bf x}^{*}={\bf 0}$ is asymptotically stable at $t=t_{0}$ if it is stable and locally attractive, i.e., $\exists\,\delta(t_{0})>0$ such that
@@
An equilibrium point which is stable but not asymptotically stable is called marginally stable, see Fig. 2. Both stability and asymptotic stability are defined at a time instant $t_{0}$. In practice, it is often desirable for a system to have a certain uniformity in its behavior. Uniform (asymptotic) stability requires that the equilibrium point is (asymptotically) stable for all $t_{0}>0$. Note that notions of uniformity are only relevant for time-varying or non-autonomous systems. For autonomous or time-invariant systems, (asymptotic) stability naturally implies uniform (asymptotic) stability. The definition of asymptotic stability does not quantify the rate of convergence. The notion of exponential stability guarantees a minimal rate of convergence. An equilibrium point ${\bf x}^{*}={\bf 0}$ is exponentially stable if $\exists\,\delta,\alpha,\lambda>0$ such that
@@
Figure 2: Geometrical implication of stability. For nonlinear systems $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t),t)$, due to the complex and rich behavior of nonlinear dynamics, various types of stability, e.g., stability, asymptotic stability, and global asymptotic stability, can be discussed. Intuitively and roughly speaking, if all solutions of the system that start out near an equilibrium point ${\bf x}^{*}$ stay near ${\bf x}^{*}$ forever, then ${\bf x}^{*}$ is stable (in the sense of Lyapunov). More strongly, if ${\bf x}^{*}$ is stable and all solutions that start out near ${\bf x}^{*}$ converge to ${\bf x}^{*}$ as $t\to\infty$, then ${\bf x}^{*}$ is asymptotically stable. ${\bf x}^{*}$ is marginally stable if it is stable but not asymptotically stable. This figure was redrawn from [236].
@@ 2.3 Lyapunov’s direct method
Table 2: Summary of Lyapunov stability theorems on equilibrium point ${\bf x}^{*}={\bf 0}$ [172].
@@ 3 Linear stability analysis
As a simple application of Lyapunov’s indirect method, linear stability analysis has been extensively applied to ecological systems, helping us better understand the intricate relationships between stability and biodiversity [11, 12, 33, 122, 150, 163, 165, 180, 263]. More significantly, under some special conditions, local stability determined by linear stability analysis implies global stability for certain ecological systems [125].
In his pioneer work, May considered an ecological system of $S$ species coexisting at a feasible equilibrium point ${\bf x}^{*}$ of a dynamical system $\dot{\bf x}(t)={\bf f}({\bf x}(t))$ that describes the time-dependent abundance vector ${\bf x}(t)$ of the $S$ species [163, 165]. Suppose that one species is subjected to a small but sudden population increase or decrease. As a result, other species populations may show immediate changes away from the equilibrium. The manipulated species itself, if with initially increased (or reduced) abundance, may begin to decline (or increase) toward the equilibrium because of self-regulation. These immediate, direct changes can be referred to as first-order effects and described by the linearized equation of the original dynamical system, i.e.,

### 07 Feb 2024 v2

# Stability of Ecological Systems: A Theoretical Review

@@ Abstract
The stability of ecological systems is a fundamental concept in ecology, which offers profound insights into species coexistence, biodiversity, and community persistence. In this article, we provide a systematic and comprehensive review on the theoretical frameworks for analyzing the stability of ecological systems. Notably, we survey various stability notions, including linear stability, sign stability, diagonal stability, D-stability, total stability, sector stability, structural stability, and higher-order stability. For each of these stability notions, we examine necessary or sufficient conditions for achieving such stability and demonstrate the intricate interplay of these conditions on the network structures of ecological systems. Finally, we explore the future prospects of these stability notions.
@@
[…]
organization=School of Data Science and Society and Department of Mathematics, University of North Carolina at Chapel Hill,city=Chapel Hill, postcode=27599, state=NC, country=USA
[…]
organization=Channing Division of Network Medicine, Department of Medicine, Brigham and Women’s Hospital, Harvard Medical School,city=Boston, postcode=02115, state=MA, country=USA
[…]
organization=Carl R. Woese Institute for Genomic Biology, Center for Artificial Intelligence and Modeling, University of Illinois at Urbana-Champaign,city=Champaign, postcode=61801, state=IL, country=USA
@@ 1 Introduction
Ecological systems are usually subject to continual disturbances or perturbations, and their responses are often characterized quantitatively as stability, a fundamental concept rooted in dynamical systems and control theory buma2020disturbances ; kefi2019advancing ; li2006global ; neubert1997alternatives ; pascual2005criticality ; saunders1975stability ; schoon2012understanding ; vsiljak1974connective ; townley2007predicting . Loosely speaking, an ecological system is stable if its state represented by a vector of its species abundances, do not change too much under small perturbations. Therefore, understanding the stability of ecological systems is particularly important due to its implications for species coexistence, biodiversity, and community persistence cottingham2001biodiversity ; goh1975stability ; grimm1992application ; harrison2001dynamical ; holling1973resilience ; levine2017beyond ; ozaki2020understanding ; tokeshi2009species . More importantly, these implications can contribute to the development of effective interventions and policies that maintain the overall health of ecological systems domptail2013managing ; francis2021management ; goh2012management ; janssen2006governing .
@@
Numerous mathematicians, physicists, and ecologists have explored the stability of ecological systems. In his pioneer work, May linearized an unspecified nonlinear dynamical system around an equilibrium to perform the local linear stability analysis of ecological systems May-Nature-1972 ; May-Book-1974 ; may1975stability . The local linear stability is completely characterized by the eigenvalues of the Jacobian matrix (of the unspecified dynamical model), which is often called the community matrix in the ecological contexts. If all the eigenvalues of the community matrix have negative real parts, the ecological system is locally stable around the equilibrium. By incorporating random matrix theory, May successfully derived a simple condition to characterize the stability of ecological systems. While there have been intense debates about the relationship between stability and complexity over the past five decades mccann2000diversity ; odenbaugh2011complex ; pimm1984complexity ; roberts1974stability ; vsiljak1975complex ; upadhyay2000stability ; veloz2020complexity ; yonatan2022complexity , May’s result offers a powerful foundation for analyzing the stability of ecological systems with diverse structural characteristics, such as interaction types Allesina-Nature-2012 , degree heterogeneity yan2017degree , self-regulation barabas2017self , modularity grilli2016modularity , and more.
@@
May’s framework does not need to know the dynamical model of the ecological system. Instead, it focuses on the Jacobian matrix of the unspecified dynamical model. Hence, the framework is model-implicit. In the past, actually many ecological models, such as the generalized Lotka-Volterra (GLV) model (1910) bhargava1989generalized ; malcai2002theoretical , Holling type II model (1959) agiza2009chaotic ; skalski2001functional , and MacArthur’s consumer-resource model (1970) advani2018statistical ; chesson1990macarthur , have been proposed to study the dynamics of ecological systems. Remarkably, the GLV model is the most commonly used model in the stability analysis of ecological systems due to its simplicity al2003stability ; altieri2021properties ; beretta1988generalization ; beretta1988global ; bunin2017ecological ; goh1977global ; kon2011age ; zeeman1998three . Most of these findings rely on Lyapunov theory, which allows us to determine the stability of a system without explicitly integrating the differential equation. Lyapunov theory encompasses two fundamental methods. One is based on linearization (e.g., May’s approach), which is often referred to as the Lyapunov indirect method, while the other is called the Lyapunov direct method. The essence of Lyapunov direct method is to construct the so-called Lyapunov function, an “energy-like” scalar function whose time variation can be viewed as the “energy dissipation.” It is not restricted to small perturbations, and in principle can be applied to all dynamical systems. However, we lack a general theory to find a suitable Lyapunov function for an arbitrary dynamical system. Instead, we have to rely on our experience and intuition. Notably, certain diagonal-type Lyapunov functions are useful for the stability analysis of ecological systems described by the GLV model Kaszkurewicz-Book-00 .
@@
The feasibility of an ecological system, which concerns whether a specific set of conditions or parameters can support the system over time, can also impact its stability aladwani2020ecological ; aparicio2023feasibility ; goh1977feasibility ; grilli2017feasibility ; paulus1995feasibility ; stone2018feasibility . Intriguingly, the relationship between feasibility and stability is noteworthy when dealing with ecological systems described by the GLV model or MacArthur’s consumer-resource model. It has been proved that under certain conditions, the feasibility of such systems can imply their stability case1979global ; stone2018feasibility . Moreover, feasibility is closely related to a concept known as structural stability, which reflects the ability of an ecological system to qualitatively maintain its dynamics and overall behavior under small disturbances or perturbations of the system model itself rohr2014structural . A great amount of metrics have been proposed to quantify the structural stability of ecological systems through feasibility analysis grilli2017feasibility ; rohr2014structural ; saavedra2014structurally .
Recently, there has been a growing interest in higher-order interactions within ecological systems, which refers to the intricate relationships that extend beyond pairwise interspecies interactions bairey2016high ; barbosa2023experimental ; gibbs2022coexistence ; grilli2017higher ; li2020advances , see Figure 1a, b. Mathematically, these higher-order interactions emerged in an ecological system can be naturally represented as a hypergraph, where its hyperedges can connect an arbitrary number of nodes berge1984hypergraphs . The resulting dynamics can be expressed in the form of polynomials chen2021controllability . The analysis of hypergraph dynamics often involves the application of tensor theory chen2020tensor ; chen2021controllability ; pickard2023hat ; surana2022hypergraph , which deals with multidimensional arrays generalized from vectors and matrices chen2022explicit ; chen2019multilinear ; chen2021multilinear . Lyapunov theory can also be leveraged to analyze the stability of polynomial dynamical systems ahmadi2019algebraic . Furthermore, the consumer-resource model can be implicitly considered as an instance of higher-order interactions, with consumer species as nodes and resources as hyperedges, see Figure 1c. Increasing evidence has revealed that higher-order interactions play a significant role in the stability of ecological systems bairey2016high ; gibbs2022coexistence ; grilli2017higher .
Table 1: Summary of stability notions that have been used to study ecological systems.
@@
Allesina-Nature-2012 ; allesina2015predicting ; barabas2017self ; baron2020dispersal ; baron2023breakdown ; galla2018dynamically ; gibbs2018effect ; gravel2016stability ; grilli2016modularity ; jacquet2016no ; krumbeck2021fluctuation ; May-Nature-1972 ; May-Book-1974 ; pigani2022delay ; poley2023generalized ; sidhom2020ecological ; stone2018feasibility ; tang2014correlation ; Yan-arXiv-14 ; yan2017degree ; yang2023time
@@
allesina2008network ; dambacher2003qualitative ; haraldsson2020model ; Jeffries-74 ; logofet1982sign ; Maybee-69 ; quirk1965qualitative ; Yamada-Networks-90
@@
arcak2011diagonal ; arcak2008passivity ; arcak2006diagonal ; 4434115 ; beretta1988generalization ; beretta1988global ; gibbs2018effect ; goh1977global ; grilli2017feasibility ; hernandez1997lotka ; hou2015global ; Kaszkurewicz-Book-00 ; kon2011age ; motee2012stability ; Redheffer-SIAM-85 ; simpson2021diagonal ; wang2012diagonal ; worz1978global
@@
arrow1958note ; chu2007equivalent ; enthoven1956theorem ; gibbs2018effect ; johnson1974sufficient ; rohr2014structural
@@
logofet2018matrices ; logofet2005stronger ; logofet1987on ; quirk1965qualitative
Sector
Stability
Generic population dynamics model (26) A semi-feasible equilibrium point is called sector stable if every solution that starts within a non-negative neighborhood remains in the same or an even larger non-negative neighborhood and eventually converges to that equilibrium point.
case1979global ; goh1978sector
@@
Andronov-1937 ; grilli2017feasibility ; Peixoto-1962 ; pettersson2020predicting ; recknagel1985analysis ; rohr2014structural ; saavedra2017structural ; saavedra2014structurally ; Smale-1961 ; Smale-1967 ; song2018structural
@@
ahmadi2019algebraic ; akimenko2021stability ; aladwani2019addition ; aparicio2023feasibility ; bairey2016high ; bieg2023stability ; butler2018stability ; case1979global ; chen2022explicit ; de2000random ; din2021discrete ; duffy2016identifying ; el2023age ; gawronski2022instability ; gibbs2022coexistence ; grilli2017higher ; kelsic2015counteraction ; schreiber1996global ; singh2021higher ; synodinos2023effects ; yoshino2008rank
Most significantly, the stability of an ecological system is profoundly determined and characterized by its underlying network structure, which captures how species influence one another and how disturbances or perturbations propagate among species. In this article, we provide a systematic and inclusive literature review on the theoretical frameworks for analyzing the stability of ecological systems. In Section 2, we briefly review the fundamentals of Lyapunov theory, which acts as a foundation for the stability analysis of ecological systems. From Section 3 to 10, we comprehensively survey various stability notions, including linear stability, sign stability, diagonal stability, D-stability, total stability, sector stability, structural stability, and higher-order stability, see Table 1. For each of these stability notions, we examine necessary or sufficient conditions that are required to establish such stability and illustrate the complicated relationships between these conditions and the network structures of ecological systems. Finally, we discuss the future prospects of these stability notions in Section 11 and conclude in Section 12.
@@ 2 Preliminaries
In this section, we review the fundamentals of Lyapunov theory, which plays a significant role in the various stability analyses of ecological systems. We adapt most of the notations and results from the comprehensive work of khalil2009lyapunov ; khalil2015nonlinear ; Murray-Book-Robotics ; sastry1999lyapunov ; Slotine-Book-91 .
@@ 2.1 Definitions of Lyapunov stability
Otherwise, ${\bf x}^{*}$ is unstable, see Figure 2. An equilibrium point ${\bf x}^{*}={\bf 0}$ is asymptotically stable at $t=t_{0}$ if it is stable and locally attractive, i.e., $\exists\,\delta(t_{0})>0$ such that
@@
An equilibrium point which is stable but not asymptotically stable is called marginally stable, see Figure 2. Both stability and asymptotic stability are defined at a time instant $t_{0}$. In practice, it is often desirable for a system to have a certain uniformity in its behavior. Uniform (asymptotic) stability requires that the equilibrium point is (asymptotically) stable for all $t_{0}>0$. Note that notions of uniformity are only relevant for time-varying or non-autonomous systems. For autonomous or time-invariant systems, (asymptotic) stability naturally implies uniform (asymptotic) stability. The definition of asymptotic stability does not quantify the rate of convergence. The notion of exponential stability guarantees a minimal rate of convergence. An equilibrium point ${\bf x}^{*}={\bf 0}$ is exponentially stable if $\exists\,\delta,\alpha,\lambda>0$ such that
@@
Figure 2: Geometrical implication of stability. For nonlinear systems $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t),t)$, due to the complex and rich behavior of nonlinear dynamics, various types of stability, e.g., stability, asymptotic stability, and global asymptotic stability, can be discussed. Intuitively and roughly speaking, if all solutions of the system that start out near an equilibrium point ${\bf x}^{*}$ stay near ${\bf x}^{*}$ forever, then ${\bf x}^{*}$ is stable (in the sense of Lyapunov). More strongly, if ${\bf x}^{*}$ is stable and all solutions that start out near ${\bf x}^{*}$ converge to ${\bf x}^{*}$ as $t\to\infty$, then ${\bf x}^{*}$ is asymptotically stable. ${\bf x}^{*}$ is marginally stable if it is stable but not asymptotically stable. This figure was redrawn from Slotine-Book-91 .
@@ 2.3 Lyapunov’s direct method
Table 2: Summary of Lyapunov stability theorems on equilibrium point ${\bf x}^{*}={\bf 0}$ Murray-Book-Robotics .
@@ 3 Linear stability analysis
As a simple application of Lyapunov’s indirect method, linear stability analysis has been extensively applied to ecological systems, helping us better understand the intricate relationships between stability and biodiversity Allesina-Nature-2012 ; allesina2015stability ; barclay1978deterministic ; hogg1989stability ; landi2018complexity ; May-Nature-1972 ; May-Book-1974 ; nunney1980stability ; weinstock1987local . More significantly, under some special conditions, local stability determined by linear stability analysis implies global stability for certain ecological systems hsu1995global .
In his pioneer work, May considered an ecological system of $S$ species coexisting at a feasible equilibrium point ${\bf x}^{*}$ of a dynamical system $\dot{\bf x}(t)={\bf f}({\bf x}(t))$ that describes the time-dependent abundance vector ${\bf x}(t)$ of the $S$ species May-Nature-1972 ; May-Book-1974 . Suppose that one species is subjected to a small but sudden population increase or decrease. As a result, other species populations may show immediate changes away from the equilibrium. The manipulated species itself, if with initially increased (or reduced) abundance, may begin to decline (or increase) toward the equilibrium because of self-regulation. These immediate, direct changes can be referred to as first-order effects and described by the linearized equation of the original dynamical system, i.e.,